Equipment and Calibration History
Page: 1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 17 18 19 20 21 22 23 24 25 26 27 28 29 30 31 32 [33] 34 35 36 37 38 39 40 41 42 43 44 45 46 47 48 49 50 51 52 53 54 55 56 57 58 59 60 61 62 63 64 65 66 67 68 69 70Ser No. | Date | Loc From | Title | Loc To | Notes |
---|---|---|---|---|---|
SI-131 7385 | 2019-04-16 | Tonzi | Removed | Lab none | Removed with the wooden cal panel for the FLIR because cows chewed the TC wires and the new cal panel is better |
LI-7500A 75H-2180 | 2019-04-23 | Lab none | Calibration | Lab none | Last calibration: 2019-01-25 AGC=102 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=24.7C, Tdew=10.9C Zero Air at 1100psi, AMP Air 408.97 at 700 Tdewspan=16.00 CO2 zero 0.9137 -> .9136 Offset: -.34ppm CO2 span 1.0262 -> 1.0266 Offset: -.2ppm H2O zero 0.8891 -> .8886 Offset: -3.0ppm H2O span 1.0199 -> 1.0236 Offset: -.07ppm Notes: |
LI-7500A 0073 | 2019-04-23 | Lab none | Calibration | Lab none | Last calibration: 2019-01-07 AGC=98.6 SDM= DAC1= DAC2= Ambient P=100.4kPa, Tair=26.0C, Tdew=11.8C Zero Air at 1100psi, AMP Air 408.97 at 700 Tdewspan=16.00 CO2 zero 0.956 -> .9563 Offset: .91ppm CO2 span 1.0081 -> 1.0094 Offset: -0.78ppm H2O zero 0.7316 -> .7336 Offset: 12.1ppm H2O span 1.0023 -> 1.0153 Offset: -.26ppm Notes: |
LI-7500A 0418 | 2019-04-22 | Lab none | Calibration | Lab none | Last calibration: 2019-01-07 AGC=95 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=24.2C, Tdew=11.4C Zero Air at 1100psi, AMP Air 408.97 at 700 Tdewspan=16.00 CO2 zero 0.8992 -> .8992 Offset: -.04ppm CO2 span 1.0033 -> 1.0011 Offset: 1.21ppm H2O zero 0.8465 -> .8464 Offset: -.4ppm H2O span 0.9832 -> .9866 Offset: -.07ppm Notes: |
LI-7500A 75H-2514 | 2019-04-22 | Lab none | Calibration | Lab none | Last calibration: 2019-01-07 AGC=102 SDM= DAC1= DAC2= Ambient P=100.6kPa, Tair=24.7C, Tdew=10.9C Zero Air at 1100psi, AMP Air 408.97 at 700 Tdewspan=16.00 CO2 zero 0.9051 -> .905 Offset: -.69ppm CO2 span 1.0167 -> 1.01 Offset: 3.39ppm H2O zero 0.8397 -> .8389 Offset: -5.5ppm H2O span 1.0177 -> 1.0155 Offset: .05ppm Notes: |
LI-7500A 75H-2669 | 2019-04-18 | Lab none | Installed | Sherman Barn Gas Analyzer CO2/H2O | Installed at SB to replace sn 75H-2180, which was brought back to the lab for calibration. |
LI-7500A 75H-2180 | 2019-04-18 | Sherman Barn Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from SB and brought back to the lab for calibration. It was replaced by sn 75H-2669. |
PA-II-SD 68c63acc07a9 | 2019-04-18 | Lab none | Installed | Bouldin Corn Air Quality Sensor | Installed again at BC after checking the new microSD card and the sensor were ok. |
CS451 70011234 | 2019-04-11 | Lab none | Installed | Bouldin Corn Water Level | Installed in a well underneath the BC tower. Sensor is 92cm below ground level. From field notes:
Top of pipe to ground = 25cm Top of pipe to sensor = 117cm Top of pipe to water = 85cm Sensor reading = 0.46 psi ~ 32cm (matches manual reading) |
PA-II-SD 68c63acc07a9 | 2019-04-11 | Bouldin Corn Air Quality Sensor | Removed | Lab none | Removed from BC and brought back to the lab for testing. The microSD card was cracked, but the sensor itself was fine. |
LI-7500A 75H-2150 | 2019-04-18 | Lab none | Installed | Sherman Wetland Gas Analyzer CO2/H2O | Installed at SW to replace sn 0418, which was brought back to the lab for calibration. |
LI-7500A 0418 | 2019-04-18 | Sherman Wetland Gas Analyzer CO2/H2O | Removed | Lab none | Removed from SW and brought back to the lab for calibration. It was replaced by sn 75H-2150. |
LI-7500A 0065 | 2019-04-18 | Lab none | Installed | East Pond Gas Analyzer CO2/H2O | Installed at EP to replace sn 0073, which was brought back to the lab for calibration. |
LI-7500A 0073 | 2019-04-18 | East Pond Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from EP and brought back to the lab for calibration. It was replaced by sn 0065. |
LI-7500A 75H-2182 | 2019-04-18 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at MB to replace sn 75H-2514, which was brought back to the lab for calibration. |
LI-7500A 75H-2514 | 2019-04-18 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab none | Removed from MB and brought back to the lab for calibration. It was replaced by sn 75H-2182. |
RF401 35785 | 2019-04-16 | Lab none | Installed | Tonzi | Install in the OSU SP1 box near the soil co2 system to connect the CR3000 to the Tower computer. Using Pakbus mode at 38400 |
WMP 1352 000022 | 2019-04-17 | Loaned | Returned | Lab none | Returned from David R. Cook at Argonne National Laboratory by way of LBL Ameriflux. Apparently David forgot who loaned them the sonic. Also it was reported that the sonic wasn't working. I had left a "serial output problem" note on its side but it seems to be spitting 10Hz binary data at 9600 baud. |
LI-7500A 0065 | 2019-04-15 | Lab none | Calibration | Lab none | Last calibration: 2019-02-11 AGC=99.2 SDM=0 DAC1=- DAC2=- Ambient P=100.16kPa, Tair=27.11C, Tdew=6.82C Zero Air 369068 at 1100psi, AMP CA04396 408.97 at 750 Tdewspan=15.99 CO2 zero 0.892 -> 0.8911 Offset: -4.4ppm CO2 span 1.0073 -> 1.0171 Offset: -5.12ppm H2O zero 0.7263 -> 0.7260 Offset: -2.4ppm H2O span 1.0104 -> 1.0118 Offset: -0.02ppm Notes: First lab calibration since factory calibration. I manually typed in the new coefficients and zero/span values. I saved a config file then opened it in a text editor and changed the calibration dates, standard gas values and WX coefficient. I saved and uploaded this config. This IRGA now has the SD* temperature coefficients and its signal strength was reset at the factory. Our zeros were lower than the factory calibration so I calibrated and saved a new lab calibration config file. |
LI-7500A 75H-2182 | 2019-04-15 | Lab none | Calibration | Lab none | Last calibration: 2018-10-24 AGC=87.6 SDM=0 DAC1=- DAC2=- Ambient P=100.19kPa, Tair=26.94C, Tdew=7.42C Zero Air 369068 at 1100psi, AMP CA04396 408.97 at 750 Tdewspan=15.99 CO2 zero 0.8844 -> 0.8842 Offset: -1.10ppm CO2 span 1.0188 -> 1.0156 Offset: 0.08ppm H2O zero 0.8449 -> 0.8459 Offset: 6.3ppm H2O span 1.0143 -> 1.0189 Offset: -0.1ppm Notes: by Joe |
LI-7500A 75H-2150 | 2019-04-15 | Lab none | Calibration | Lab none | Last calibration: 2018-10-24 AGC=97.2 SDM=0 DAC1=- DAC2=- Ambient P=100.19kPa, Tair=26.84C, Tdew=7.32C Zero Air 369068 at 1100psi, AMP CA04396 408.97 at 750 Tdewspan=15.99 CO2 zero 0.9021 -> 0.9014 Offset: -3.16ppm CO2 span 1.005 -> 1.0030 Offset: 0.86ppm H2O zero 0.866 -> 0.8671 Offset: 8.0ppm H2O span 1.0069 -> 1.0077 Offset: -0.02ppm Notes: By Joe |
LI-7500A 75H-2669 | 2019-04-15 | Lab none | Calibration | Lab none | Last calibration: 2018-10-10 AGC=103.8 SDM=0 DAC1=- DAC2=- Ambient P=100.16kPa, Tair=24.85C, Tdew=6.62C Zero Air 369068 at 1150psi, AMP CA04396 408.97 at 800 Tdewspan=15.99 CO2 zero 0.9188 -> 0.9186 Offset: -1.10ppm CO2 span 1.0106 -> 1.0134 Offset: -1.52ppm H2O zero 0.8992 -> 0.8991 Offset: -0.3ppm H2O span 1.0034 -> 1.0017 Offset: 0.03ppm Notes: Very stable - Joe |
LI-7700 TG1-0421 | 2019-04-11 | Bouldin Alfalfa Gas Analyzer CH4 | Upgrade | Bouldin Alfalfa Gas Analyzer CH4 | Upgraded LI-7700 sn TG1-0421 firmware from 1.0.23 to 1.0.27. To connect, Joe had to change the laptop network subnet from 192.168.2.5 to 192.168.1.5 (to match the LI-7700 IP address of 192.168.1.77), even though the subnet mask was 255.255.255.0. We then restarted the instrument and calibrated LI-7700 sn TG1-0421. |
USB2000-VIS-NIR USB2G4461 | 2016-10-11 | East End | Calibration | East End | Calibration Location: East End Last calibration: 2016-04-13 RSSI=82%, Optical RH=21% Ambient P=101.6kPa, Tair=21.1C zero air at 800psi, AMP CH4 4.03 at 1400 CH4 Zero Offset: 0.02ppm CH4 Span Offset: 0.01ppm Notes: Breezy |
LI-7700 TG1-0357 | 2019-04-11 | Sherman Barn Gas Analyzer CH4 | Calibration | Sherman Barn Gas Analyzer CH4 | Calibration Location: Sherman Barn Last calibration: 2018-07-10 RSSI=70%, Optical RH=3% Ambient P=101.9kPa, Tair=16.5C Zero Air M4002 at 1800psi, AMP CA03923 3.121 at 1800 CH4 Zero Offset: 0.035ppm CH4 Span Offset: -0.002ppm Notes: Windy 3m/s to 7.5m/s |
LI-7700 TG1-0211 | 2019-04-11 | East Pond Gas Analyzer CH4 | Calibration | East Pond Gas Analyzer CH4 | Calibration Location: East Pond Last calibration: 2018-10-26 RSSI=70%, Optical RH=14% Ambient P=101.9kPa, Tair=17.1C Zero Air M4002 at 1600psi, AMP CA03923 3.121 at 1800 CH4 Zero Offset: 0.014ppm CH4 Span Offset: 0.003ppm Notes: Windy day |
LI-7700 TG1-0155 | 2019-04-11 | Bouldin Corn Gas Analyzer CH4 | Calibration | Bouldin Corn Gas Analyzer CH4 | Calibration Location: Bouldin Corn Last calibration: 2018-10-25 RSSI=87%, Optical RH=22% Ambient P=101.5kPa, Tair=19.1C Zero Air M4002 at 1600psi, AMP C03923 3.121 at 1600 CH4 Zero Offset: 0.024ppm CH4 Span Offset: -0.007ppm Notes: Windy |
LI-7700 TG1-0424 | 2019-04-11 | Sherman Wetland Gas Analyzer CH4 | Calibration | Sherman Wetland Gas Analyzer CH4 | Calibration Location: Sherman Wetland
Last calibration: 2018-10-24 RSSI=77%, Optical RH=13% Ambient P=102.0kPa, Tair=15.4C Zero Air M4002 at 1800psi, AMP CA03923 3.121 at 1900 CH4 Zero Offset: 0.051ppm CH4 Span Offset: 0.024ppm Notes: Windy day 3m/s to 7m/s |
LI-7700 TG1-0423 | 2019-04-11 | East End Gas Analyzer CH4 | Calibration | East End Gas Analyzer CH4 | Calibration Location: East End Last calibration: 2018-10-30 RSSI=73%, Optical RH=11% Ambient P=101.9kPa, Tair=16.5C Zero Air M4002 at 1700psi, AMP CA03923 3.121 at 1800 CH4 Zero Offset: 0.036ppm CH4 Span Offset: 0.013ppm Notes: Windy day |
LI-7700 TG1-0421 | 2019-04-11 | Bouldin Alfalfa Gas Analyzer CH4 | Calibration | Bouldin Alfalfa Gas Analyzer CH4 | Calibration Location: Bouldin Alfalfa Last calibration: 2018-11-06 RSSI=83%, Optical RH=17% Ambient P=101.3kPa, Tair=19.2C Zero Air M4002 at 1600psi, AMP CA03923 3.121 at 1600 CH4 Zero Offset: 0.037ppm CH4 Span Offset: 0.007ppm Notes: Breezy day, upgraded firmware from v1.0.23 to v1.0.27 before calibration |
LI-7700 TG1-0224 | 2019-04-11 | West Pond Gas Analyzer CH4 | Calibration | West Pond Gas Analyzer CH4 | Calibration Location: West Pond Last calibration: 2018-10-25 RSSI=71%, Optical RH=24% Ambient P=101.8kPa, Tair=17.2C Zero Air M4002 at 1600psi, AMP CA03923 3.1210 at 1700 CH4 Zero Offset: -0.005ppm CH4 Span Offset: 0.008ppm Notes: Windy day. Optical RH high - needs new desiccant. |
PRB-K Type #8 00000128B7B90 | 2019-03-29 | Vaira Moisture Point Gate | Removed | Lab none | Removed from the location near the gate. This probe has two broken pins top and bottom and is no longer making readings. |
LI-6400 PSC-0785 | 2019-03-25 | Lab none | Calibration | Lab none | Camilo Rey-Sanchez and Daphne Szutu
Camilo cleaned the light-source leaf chamber (it wasn't too dusty) and attached it to this console for calibration. We connected the reference and sample inlet to the cal gases with a y-junction. Camilo had checked the zeros using the chemicals last week and found them unsatisfactory (we didn't write them down). We set the zero and the span using cal gas and dewpoint generator. Zero: Cal Gas (zero air), flow at 1SCFH: Before: CO2_S:~-15 CO2_R: ~-15 H2O_S: ~-2 H2O_R: ~-2 After: *All fluctuating around zero. Span with AMP 408.97ppm cal gas, flow at 1SCFH: Before: CO2_R: 413 CO2_S: 423 After: *Fluctuating around 409 Span with 16.0C dew point (2019-03-26), flow at 500 CC/min: Before: H2O_R: 17.50 C H2O_S: 18.52 C After: H2O_R: 16.00 C H2O_S: 16.00 C |
LI-6400 PSC-2474 | 2019-03-25 | Lab none | Calibration | Lab none | Camilo Rey-Sanchez and Daphne Szutu
Camilo attached a leaf chamber with a transparent top. He cleaned the chamber, but it wasn't as dusty as he suspected. I changed the time from 15:26 to 13:38. We connected the reference and sample inlet to the cal gases/dewpoint generator with a y-junction. We checked the zero using the chemicals and they seemed ok, but we decided to calibrate zero/spans using the the cal gases and dewpoint generator anyways. Zero: Checked using chemicals: CO2_S: 1.4 CO2_R: -0.1 H2O_S: 0.55 H2O_R: 1.4 Set zero using cal gas: Cal gas = zero air, 1100 psi, 1 SCFH flow Before: CO2_S:-0.9 CO2_R: 4.2 H2O_S: -0.043 H2O_R: -0.592 After: *All fluctuating around zero. Span with 408.97ppm cal gas from AMP: Cal gas = AMP cal gas at 408.97ppm, 800 psi, 1 SCFH flow Before: CO2_R: 408.7 CO2_S: 407.3 After: CO2_R: Did not adjust CO2_S: 409.0 Span with 16.0C dew point: Before: H2O_R: ?? C H2O_S:?? C After: ?? |
CR1000 66252 | 2019-03-14 | Bouldin Alfalfa Data Logger Met | Maintenance | Bouldin Alfalfa Data Logger Met | The CR1000 packbus id was change from 1 to 52 |
2300-000 2300565 | 2019-02-22 | Mayberry Fluorometer | Removed | Lab none | Removed from MB and brought back to the lab to swap out its brushes and other maintenace. |
LI-7500A 0065 | 2019-02-11 | Lab none | Calibration | Lab none | Last calibration: 2018-09-26 AGC= SDM= DAC1= DAC2= Ambient P=kPa, Tair=C, Tdew=C at psi, at Tdewspan=12.00 CO2 zero 0.8903 -> .8920 Offset: ppm CO2 span 1.0184 -> 1.0073 Offset: ppm H2O zero 0.7086 -> 07263 Offset: ppm H2O span 1.0155 -> 1.0104 Offset: ppm Notes: CX=34338.1, WX=34091.7. Calibrated at Licor factory after source/chopper replacement. |
LI-7500A 0065 | 2019-02-22 | Factory | Received | Lab none | Received from Licor after repair and calibration. Service notes: Inspected the unit and found the source lamp intermittent. Replaced the defective source and the chopper motor per the customer's request. Subjected the unit to environmental tests and calibrated. |
LI-7500A 0065 | 2019-01-21 | Lab none | Shipped | Factory | Shipped back to factory for repair. Received by Licor on 2019-01-21. |
SRS-Ni NDVI 1234903795 | 2019-02-20 | Lab none | Installed | Vaira Narrow Band Reflected | Installed on the Vaira Tower in the reflected position. |
SRS-Ni NDVI 1334903221 | 2019-02-20 | Lab none | Installed | Vaira Narrow Band Incoming | Installed on the Vaira Tower in the incoming position |
PA-II-SD 68c63acc07a9 | 2019-02-12 | Lab none | Installed | Bouldin Corn Air Quality Sensor | Installed at BC. It is bolted to the bottom side of the upper boardwalk. |
LI-7700 TG1-0421 | 2019-02-12 | Bouldin Alfalfa Gas Analyzer CH4 | Tested | Bouldin Alfalfa Gas Analyzer CH4 | Tried to upgrade firmware from 1.0.23 to 1.0.27, but there was an error, "Fatal: Network Error: Connection Timed out." |
LI-7700 TG1-0424 | 2019-02-12 | Sherman Wetland Gas Analyzer CH4 | Upgrade | Sherman Wetland Gas Analyzer CH4 | Upgraded firmware (embedded software) from 1.0.23 to 1.0.27. Licor change log: "Expanded a buffer in the DSP to prevent phase changes in the DSP data caused when the buffer filled up. This fixes the synchronization loss and resultant data loss between the LI-7550 and LI-7700." |
LI-7700 TG1-0423 | 2019-02-12 | East End Gas Analyzer CH4 | Upgrade | East End Gas Analyzer CH4 | Upgraded firmware (embedded software) from 1.0.23 to 1.0.27. Licor change log: "Expanded a buffer in the DSP to prevent phase changes in the DSP data caused when the buffer filled up. This fixes the synchronization loss and resultant data loss between the LI-7550 and LI-7700." |
LI-7500A 75H-2667 | 2019-02-12 | Lab none | Installed | East End Gas Analyzer CO2/H2O | Installed at EE to replace sn 75H-2150, which was brought back to the lab for calibration. |
LI-7500A 75H-2150 | 2019-02-12 | East End Gas Analyzer CO2/H2O | Removed | Lab none | Removed from EE and brought back to the lab for calibration. It was replaced by sn 75H-2667. |
LI-7500A 75H-2668 | 2019-02-12 | Lab none | Installed | West Pond Gas Analyzer CO2/H2O | Installed at WP to replace 75H-2182, which was brought back to the lab for calibration. |
LI-7500A 75H-2182 | 2019-02-12 | West Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from WP and brought back to the lab for calibration. It was replaced by 75H-2668. |
LI-7500A 75H-2668 | 2019-02-11 | Lab none | Calibration | Lab none | Last calibration: 2018-10-08 AGC=99 SDM=0 DAC1=0 DAC2=0 Ambient P=101.1kPa, Tair=22.5C, Tdew=3.7C Zero Air at 1150psi, AMP Air 408.97 at 800 Tdewspan=16.00 CO2 zero 0.9394 -> .939 Offset: -1.51ppm CO2 span 1.0164 -> .018 Offset: -.81ppm H2O zero 0.9057 -> .9073 Offset: 8.8ppm H2O span 1.0157 -> 1.0117 Offset: 0.08ppm Notes: Dewpoint generator dewpoint kept falling b/c it was low on water. I added more DI water and it seemed fine. |
LI-7500A 75H-2667 | 2019-02-11 | Lab none | Calibration | Lab none | Last calibration: 2018-10-10 AGC=104 SDM=0 DAC1=0 DAC2=0 Ambient P=101.2kPa, Tair=25.1C, Tdew=2.5C Zero Air at 1175psi, AMP Air 408.97 at 800 Tdewspan=16.00 CO2 zero 0.9412 -> .9425 Offset: 4.87ppm CO2 span 1.0145 -> 1.0143 Offset: 0.15ppm H2O zero 0.8367 -> .8349 Offset: -11.1ppm H2O span 1.0218 -> 1.0222 Offset: -.01ppm Notes: Spanned at "15.99C" in 7550 software with dewpt generator set at 16.00. |
WM 1590 120907 | 2019-02-06 | Loaned | Returned | Lab none | Returned by Tibisay Perez who used it at Whendee Silver's compost pile experiment in Marin |
WM 1590 153805 | 2019-02-06 | Loaned | Returned | Lab none | Returned by Tibisay Perez who used it at Whendee Silver's compost pile experiment in Marin |
LI-7500A 75H-2180 | 2019-01-31 | Lab none | Installed | Sherman Barn Gas Analyzer CO2/H2O | Installed at SB to replace sn 75H-2669, which was brought back to the lab for calibration. |
LI-7500A 75H-2669 | 2019-01-31 | Sherman Barn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from SB and brought back to the lab for calibration. It was replaced by 75H-2180. |
LI-7500A 1594 | 2019-01-31 | Lab none | Installed | Bouldin Corn Gas Analyzer CO2/H2O | Installed at BC to replace sn 75H-2668, which was brought back to the lab for calibration. |
LI-7500A 75H-2668 | 2019-01-31 | Bouldin Corn Gas Analyzer CO2/H2O | Removed | Lab none | Removed from BC and brought back to the lab for calibration. It was replaced by sn 1594. |
LI-7500A 0035 | 2019-01-31 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CO2/H2O | Installed at BA to replace sn 75H-2667, which was brought back to the lab for calibration. |
LI-7500A 75H-2667 | 2019-01-31 | Bouldin Alfalfa Gas Analyzer CO2/H2O | Removed | Lab none | Removed from BA and brought back to the lab for calibration. It was replaced by sn 0035. |
CR1000 50889 | 2019-01-31 | East Pond Data Logger | Serviced | East Pond Data Logger | Changed Pakbus address from 1 to 89. |
HMP60 L1740420 | 2019-01-31 | Lab none | Installed | East End RH & Temp | Installed at EE to replace HMP45 sn U3030046, which was removed for bad RH readings. |
HMP45AC U3030046 | 2019-01-31 | East End RH & Temp | Removed | Lab none | Removed from EE b/c RH was not working (ranged only b/t 10 and 13%). It was replaced by HMP60 sn L1740420. |
CR1000 47109 | 2019-01-31 | West Pond Data Logger Met | Serviced | West Pond Data Logger Met | Changed Pakbus address from 1 to 9. |
Mark I A000251 | 2017-09-07 | Lab none | Installed | Twitchell Alfalfa Weather Station | Installed at Twitchell Alfalfa with Joe Berry. This was just a quick stop on our way to Tonzi and Vaira. |
Line Quantum LQ1570 | 2019-01-29 | Factory | Disposed | Disposed | Disposed of at factory. Licor determined it was necessary to replace practically the whole innards of the sensor. |
Line Quantum LQ1570 | 2019-01-16 | Lab none | Shipped | Factory | Joe had cleaned up and tested this sensor, but we saw only noise in the data, so we returned to Licor for inspection and repair, RMA 211671. |
LI-7500A 1594 | 2019-01-30 | Lab none | Calibration | Lab none | Last calibration: 2018-10-03 AGC=98.88 SDM=0 DAC1=0 DAC2=0 Ambient P=100.07kPa, Tair=27.07C, Tdew=10.85C Zero Air at 1200psi, AMP Air 408.97 at 825 Tdewspan=16.00 CO2 zero 0.8644 -> .8643 Offset: -.16ppm CO2 span 1.015 -> 1.0157 Offset: .5ppm H2O zero 0.887 -> .8891 Offset: 11.9ppm H2O span 1.0056 -> 1.0043 Offset: .02ppm Notes: Calibrated by Jiangong. |
LI-7500A 0035 | 2019-01-28 | Lab none | Calibration | Lab none | Last calibration: 2018-10-02 AGC=99.2 SDM=0 DAC1=0 DAC2=0 Ambient P=100.7kPa, Tair=24.0C, Tdew=10.04C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9375 -> .9375 Offset: .41ppm CO2 span 1.0142 -> 1.0193 Offset: 2.5ppm H2O zero 0.8981 -> .8969 Offset: -7.6ppm H2O span 1.0277 -> 1.0023 Offset: .54ppm Notes: Calibrated by Jiangong. Recorded on the from for sn 0035 but the correct config file (sn 0065) was loaded before the calibration, so the calibration is still good. |
LI-7500A 75H-2180 | 2019-01-25 | Lab none | Calibration | Lab none | Last calibration: 2018-10-02 AGC=98 SDM=0 DAC1=0 DAC2=0 Ambient P=101.2kPa, Tair=26.9C, Tdew=5.3C Zero Air at 1200psi, AMP Air 408.97 at 875 Tdewspan=16.00 CO2 zero 0.9137 -> .9137 Offset: 407.3ppm CO2 span 1.023 -> 1.0262 Offset: -1.67ppm H2O zero 0.883 -> .8891 Offset: 38.5ppm H2O span 0.9893 -> 1.0199 Offset: -.63ppm Notes: Date of last zeros were 2018-06-25, but the existing coefficients matched database entry from 2018-10-02. Training Jiangong. |
CR1000 47108 | 2019-01-24 | Sherman Barn Data Logger Met | Serviced | Sherman Barn Data Logger Met | Changed Pakbus address from 1 to 8. |
CR1000 38211 | 2019-01-24 | Sherman Wetland Data Logger Met | Serviced | Sherman Wetland Data Logger Met | Changed Pakbus address from 1 to 11. |
HMP45AC W0730036 | 2019-01-24 | Lab none | Installed | Sherman Wetland RH & Temp | Installed at SW to replace T2130005, which was brought back to the lab for inspection. |
HMP45AC T2130005 | 2019-01-24 | Sherman Wetland RH & Temp | Removed | Lab none | Brought back to the lab from SW for inspection b/c of data drops/spikes. It was replaced by sn W0730036. |
CR1000 47110 | 2019-01-24 | East End Data Logger Met | Serviced | East End Data Logger Met | Changed Pakbus address from 1 to 12. |
HMP45AC U3030046 | 2019-01-24 | Lab none | Installed | East End RH & Temp | Installed at EE to replace G4540023, which was brought back to the lab for inspection. |
HMP60 G4540023 | 2019-01-24 | East End RH & Temp | Removed | Lab none | Brought back to the lab from EE for inspection b/c of data drops/spikes. It was replaced by HMP45 sn U3030046. |
CR1000 38212 | 2019-01-24 | Mayberry Data Logger Tule | Serviced | Mayberry Data Logger Tule | Changed Pakbus address from 1 to 12. |
HMP45A U2120013 | 2019-01-24 | Mayberry RH & Temp | Serviced | Mayberry RH & Temp | Short in the 12V wire at the splice point. Repaired in the field with a screw terminal, but should be brought back to the lab for a full repair. |
HMP45AC W0730036 | 2019-01-23 | Lab none | Maintenance | Lab none | Spliced on a longer cable so total length is ~20'. New wiring colors are: Orange-Temperature, Blue-RH, White-SignalReference, Red-12V, Black-G, Bare/Shield-AG |
HMP45AC U3030046 | 2019-01-23 | Lab none | Maintenance | Lab none | Spliced on a longer cable so total length is ~24'. New wiring colors are: Orange-Temperature, Blue-RH, White-SignalReference, Green-Control, Red-12V, Black-G, Bare/Shield-AG |
CR10X 20138 | 2019-01-13 | Lab none | Installed | Tonzi | Installed near the Floor solar panels to record temperatures from a concrete cal panel for the FLIR thermal camera. |
LI-7500A 0073 | 2019-01-18 | Lab none | Installed | East Pond Gas Analyzer CO2/H2O | Installed at EP to replace sn 0035, which was brought back to the lab for calibration. |
LI-7500A 0035 | 2019-01-18 | East Pond Gas Analyzer CO2/H2O | Removed | Lab none | Removed from EP and brought back to the lab for calibration. it was replaced by sn 0073. |
LI-7500A 0418 | 2019-01-18 | Lab none | Installed | Sherman Wetland Gas Analyzer CO2/H2O | Installed at SW to replace 75H-2180, which was brought back to the lab for calibration. |
LI-7500A 75H-2180 | 2019-01-18 | Sherman Wetland Gas Analyzer CO2/H2O | Removed | Lab none | Removed from SW and brought back to the lab for calibration. It was replaced by sn 0418. |
LI-7500A 75H-2514 | 2019-01-18 | Lab none | Installed | Mayberry Gas Analyzer CO2/H2O | Installed at MB to replace sn 1594, which was brought back to the lab for calibration. |
LI-7500A 1594 | 2019-01-18 | Mayberry Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from MB and brought back to the lab for calibration. It was replaced by sn 75H-2514. |
LI-7500A 0073 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-09-26 AGC=100 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=18.7C, Tdew=9.5C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9569 -> .9560 Offset: -2.67ppm CO2 span 1.0112 -> 1.0081 Offset: 1.65ppm H2O zero 0.7314 -> .7316 Offset: .6ppm H2O span 1.0183 -> 1.0023 Offset: .35ppm Notes: |
LI-7500A 0418 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-10-10 AGC=97.2 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=17.9C, Tdew=9.7C Zero Air at 1200psi, AMP Air 408.97 at 900 Tdewspan=16.00 CO2 zero 0.8982 -> .8992 Offset: 3.69ppm CO2 span 0.9965 -> 1.0033 Offset: -3.58ppm H2O zero 0.8454 -> .8465 Offset: 6.1ppm H2O span 0.9813 -> .9832 Offset: -.03ppm Notes: |
LI-7500A 75H-2514 | 2019-01-07 | Lab none | Calibration | Lab none | Last calibration: 2018-09-26 AGC=102.5 SDM=0 DAC1=0 DAC2=0 Ambient P=100.9kPa, Tair=18.0C, Tdew=10.1C Zero Air at 1200psi, AMP Air 408.97 at 850 Tdewspan=16.00 CO2 zero 0.9045 -> .9051 Offset: 2.36ppm CO2 span 1.0036 -> 1.0167 Offset: -6.83ppm H2O zero 0.837 -> .8397 Offset: 19.5ppm H2O span 1.0165 -> 1.0177 Offset: -.03ppm Notes: |
LI-7700 TG1-0421 | 2018-11-07 | Lab none | Installed | Bouldin Alfalfa Gas Analyzer CH4 | Installed at BA. The previous LI-7700 sn TG1-0424 was removed for a power board swap. |
LI-7500A 0041 | 2019-01-04 | Lab none | Installed | Tonzi-Floor | Installed on the Tonzi Floor to replace sn 75H-2514 for calibration. |
LI-7500A 75H-2514 | 2019-01-04 | Tonzi-Floor Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Tonzi Floor for calibration. Replaced by sn 0041 |
LI-7500A 0042 | 2019-01-04 | Lab none | Installed | Tonzi-Tower Gas Analyzer CO2/H2O | Installed on the Tonzi Tower to replace sn 0073 for calibration. |
LI-7500A 0073 | 2019-01-04 | Tonzi-Tower Gas Analyzer CO2/H2O | Removed | Lab Gas Analyzer CO2/H2O | Removed from the Tonzi Tower for calibration. Replaced by sn 0042 |
RF400 21519 | 2019-01-04 | Vaira | Removed | Lab none | Remove from the pond data logger box for testing in the lab. The antenna cable was broken. The radio has not been used in the field for a long while. |
RF400 21520 | 2019-01-04 | Vaira | Removed | Lab none | Remove from the met data logger box for testing in the lab. The radio has not been used in the field for a long while. |
LI-7500A 75H-2176 | 2019-01-04 | Lab none | Installed | Vaira Gas Analyzer CO2/H2O | Installed on the Vaira tower to replace sn 0418 for calibration. |
LI-7500A 0418 | 2019-01-04 | Vaira Gas Analyzer CO2/H2O | Removed | Lab none | Removed from the Vaira tower for calibration. Replaced by sn 75H-2176. |
GP-001 GP20140019 | 2019-01-03 | East End CO2 probe | Removed | Lab none | Removed and brought back to the lab for testing. |